15x^2+31x+2 what is the factor of this equation
Found 2 solutions by ewatrrr, TimothyLamb:
Answer by ewatrrr(24785)

15x^2+31x+2 = 0
(15x + 1)(x + 2) = 0)
F First terms
O Outside terms
I Inside terms
L Last terms

Answer by TimothyLamb(4379)

15x^2 + 31x + 2 = 0
---
the above quadratic equation is in standard form, with a=15, b=31 and c=2
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
15 31 2
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = -0.066666666 = -2/30
x = -2
---
by factoring:
---
15xx + 31x + 2 = 0
(15xx + 31x + 2 = 0)/15
xx + (31/15)x + 2/15 = 0
xx + (2/2)(31/15)x + (2/2)(2/15) = 0
xx + (62/30)x + (4/30) = 0
---
factoring:
(x + 2/30)(x + 60/30) = 0
---
check the above factoring for correctness:
xx + (60/30)x + (2/30)x + (2/30)(60/30) = 0
xx + (62/30)x + (2*60/30*30) = 0
xx + (62/30)x + (2*2/30) = 0
xx + (62/30)x + (4/30) = 0
---
simplify the factoring:
(x + 2/30)(x + 2) = 0
---
two real roots:
x = -2/30
x = -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php